OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through (2, -1) and parallel to a line with slope of 3/4

OpenStudy (anonymous):

y= mx+ c since the line is parallel to the line with slope(gradient), m= 3/4 sub in (2, -1) -1= 3/4(2)+ c c= -5/2 so y= 3/4x- 5/2

OpenStudy (anonymous):

How can I change the fractions into a number?

OpenStudy (anonymous):

-5/2= -2.5 3/4= 0.75 insert the numbers into the calculator?

OpenStudy (anonymous):

I can't insert decimals..

OpenStudy (anonymous):

@jocelynong can you help me :(

OpenStudy (anonymous):

you need help with the question or.. ?

OpenStudy (anonymous):

The question. I can't insert any decimals or fractions for the answers,,so is there a way to get an answer without any decimals and fractions